OpenStudy (anonymous):

can some one please answer my question fast please ]ills dog fiona weighs 8 1/4 pounds her dog felix weighs 1 1/2 times as much as fiona how much does felix weigh can some one please answer my question fast please ]ills dog fiona weighs 8 1/4 pounds her dog felix weighs 1 1/2 times as much as fiona how much does felix weigh @Mathematics

OpenStudy (anonymous):

\[8\tfrac{1}{4}\times 1\tfrac{1}{2}\] is what you need to compute

OpenStudy (anonymous):

can u do it 4 me please and show me how u got it

OpenStudy (anonymous):

start with "improper fractions" get \[\frac{33}{4}\times \frac{3}{2}\] the multiply straight across, then turn back to a mixed number

OpenStudy (anonymous):

do you know how to change \[8\tfrac{1}{4}\] into \[\frac{33}{4}\]?

OpenStudy (anonymous):

i dont know how to do dat

OpenStudy (anonymous):

\[8\tfrac{1}{4}=\frac{4\times 8+1}{4}=\frac{32+1}{4}=\frac{33}{4}\] is the method

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

no that is how you make a mixed number into an improper fraction you find your answer from computing \[\frac{33}{4}\times \frac{3}{2}\]

OpenStudy (anonymous):

multiply straight across top and bottom

OpenStudy (anonymous):

can u just please show me how u did it and give me the answer the i will give u meadal

OpenStudy (anonymous):

@meekmillz23 - how is it going to help you if all you get are steps and answer? I think Satellite73 is trying to teach you how to solve such problems. You should learn so you can do such problems by yourself in the future.

OpenStudy (anonymous):

i just go here to get quick answers and i ask them to show me how me how i got it because i will practice how they did it can u answer my other question

OpenStudy (anonymous):

As Satellite73 already did most of the steps and showed you that all you need to do is multiply (33/4) and (3/2). Can you do that?

OpenStudy (anonymous):

no how did he get 33/4 x 3/2 i come here to get the answer beacause my teacher is very strick

OpenStudy (anonymous):

8 1/4 = 33/4 Because 8 1/4 = 8 + 1/4 = (8*4 + 1)/4 = (32+1)/4 = 33/4 1 1/2 = 3/2 Because 1 1/2 = 1 + 1/2 = (1*2+1)/2 = (2+1)/2 = 3/2

OpenStudy (anonymous):

what operation is per,how many

OpenStudy (anonymous):

divided by

OpenStudy (anonymous):

miles per hour \[\text{miles}\div{\text{hours}}\] for example

OpenStudy (anonymous):

so per is dived by and how much is wat